# CLiC-it 2024: Tenth Italian Conference on Computational Linguistics, Dec 04 — 06, 2024, Pisa, Italy

### v1

@@ Abstract
Questo lavoro chiede se i modelli linguistici multilingue preaddestrati catturino rappresentazioni linguistiche astratte valide attraverso svariate lingue. In nostro approccio sviluppa dati sintetici curati su larga scala, con proprietà specifiche, e li utilizza per studiare le rappresentazioni di frasi costruite con modelli linguistici preaddestrati. Utilizziamo un nuovo task a scelta multipla e i dati afferenti, le Blackbird Language Matrices (BLM), per concentrarci su uno specifico fenomeno strutturale grammaticale - l’accordo tra il soggetto e il verbo - in diverse lingue. Per trovare la soluzione corretta a questo task è necessario un sistema che rilevi modelli e paradigmi linguistici complessi nelle rappresentazioni testuali. Utilizzando un’architettura a due livelli che risolve il problema in due fasi - prima impara gli oggetti sintattici e le loro proprietà nelle singole frasi e poi ne ricava gli elementi comuni - dimostriamo che, nonostante siano stati addestrati su testi multilingue in modo coerente, i modelli linguistici multilingue preaddestrati presentano differenze specifiche per ogni lingua e inoltre la struttura sintattica non è condivisa, nemmeno tra lingue tipologicamente molto vicine.
@@ 1 Introduction
To investigate this question, we use a seemingly simple task – subject-verb agreement. Subject-verb agreement is often used to test the syntactic abilities of deep neural networks [7, 8, 9, 10], because while apparently simple and linear, it is in fact structurally, and theoretically, complex, and requires computing the structural distance between the subject and the verb. It has an added useful dimension – it relies on syntactic structure and grammatical number information that many languages share.
In previous work we have shown that simple structural information – the chunk structure of a sentence – which can be leveraged to determine subject-verb agreement, or to contribute towards more semantic tasks, can be detected in the sentence embeddings obtained from a pre-trained model [11]. This does not cast light on whether the discovered structure is deeper and more abstract, or it is rather just a reflection of surface indicators, such as function words or morphological markers.
To test this idea, we set up an experiment covering four languages: English, French, Italian and Romanian. These languages, while different, have shared properties that make sharing of syntactic structure a reasonable expectation if the pretrained multilingual model does indeed discover and encode syntactic structure. We use parallel datasets in the four languages, built by (approximately) translating the BLM-AgrF dataset [12], a multiple-choice linguistic test inspired from the Raven Progressive Matrices visual intelligence test, previously used to explore subject-verb agreement in French.
Our cross-lingual and multilingual experiments show poor transfer across languages, even those most related, like Italian and French. This indicates that the manner in which pretrained models encode syntactic information is based on shallow and language-specific clues, from which they are not yet able to take the step towards abstracting grammatical structure.
Our work provides two contributions: (i) four parallel datasets – on English, French, Italian and Romanian, focused on subject-verb agreement; (ii) cross-lingual testing of a multilingual pretrained model, to explore the degree to which syntactic structure information is shared across different languages.
@@ 2 BLM task and BLM-Agr datasets
Raven’s progressive matrices (RPMs) are IQ tests consisting of a sequence of images, called the context, connected in a logical sequence by underlying generative rules [13]. The task is to determine the missing element in this visual sequence, the answer. The candidate answers are constructed to be similar enough that the solution can be found only if the rules are identified correctly.
@@
1 NP-sing PP1-sing VP-sing
2 NP-plur PP1-sing VP-plur
3 NP-sing PP1-plur VP-sing
4 NP-plur PP1-plur VP-plur
5 NP-sing PP1-sing PP2-sing VP-sing
6 NP-plur PP1-sing PP2-sing VP-plur
7 NP-sing PP1-plur PP2-sing VP-sing
@@
Answer
1 NP-sing PP1-sing et NP2 VP-sing Coord
2 NP-plur PP1-plur PP2-sing VP-plur correct
3 NP-sing PP-sing VP-sing WNA
4 NP-plur PP1-sing PP1-sing VP-plur WN1
5 NP-plur PP1-plur PP2-plur VP-plur WN2
6 NP-sing PP1-sing PP2-sing VP-plur AEV
7 NP-sing PP1-plur PP2-sing VP-plur AEN1
8 NP-sing PP1-sing PP2-plur VP-plur AEN2
Figure 1: BLM instances for verb-subject agreement, with two attractors. WNA= wrong number of attractors; WN1= wrong nr. for 1st attractor noun (N1); WN2= wrong nr. for 2nd attractor noun (N2); AEV=agreement error on the verb; AEN1=agreement error on N1; AEN2=agreement error on N2.
@@ 8 NP-sing PP1-sing PP2-plur VP-plur AEN2
Such an approach is be very useful for probing language models, as it allows to test whether they indeed detect the relevant linguistic objects and their properties, and whether they use this information to find larger patterns. We developed such a framework, called Blackbird Language Matrices (BLMs) [15] and developed several datasets [12, 16]. Figure 1 illustrates the template of a BLM agreement matrix, with the different linguistic objects – chunks/phrases – and their relevant properties, in this case grammatical number. Examples in all languages under investigation are provided in Appendix B.
@@
A BLM problem for subject-verb agreement consists of a context set of seven sentences that share the subject-verb agreement phenomenon, but differ in other aspects – e.g. number of linearly intervening noun phrases between the subject and the verb, called attractors because they can interfere with the agreement, different grammatical numbers for these attractors, and different clause structures. The sequence is generated by a rule of progression of number of attractors, and alternation in the grammatical number of the different phrases. Each context is paired with a set of candidate answers generated from the correct answer by altering it to produce minimally contrastive error types. They help investigate the kind of information and structure learned, by error analysis.
The seed data for French was created by completing data from [17]. From this we generated a dataset that comprise three subsets of increasing lexical complexity (details in [12]): Types I, II, III, corresponding to different amounts of lexical variation within a problem instance. Each subset contains three clause structures (main and embedded with prepositional phrases and main with relative clause), uniformly distributed within the data.
The datasets in English, Italian and Romanian were created by manually translating the seed French sentences into the other languages by native (Italian and Romanian) and near-native (English) speakers. The internal structure in these languages is very similar so translations are approximately parallel. The differences lie in the treatment of preposition and determiner sequences that must be conflated into one word in some cases in Italian and French, but not in English. French and Italian use number-specific determiners and inflections, while Romanian and English encode grammatical number exclusively through inflections. In English most plural forms are marked by a suffix. Romanian has more variation, and noun inflections also encode case. Determiners are separate tokens, which are overt indicators of grammatical number and of phrase boundaries, whereas inflections may or may not be tokenized separately.
@@
The dataset consists of triples $(in,out^{+},out^{-})$, where $in$ is an input sentence, $out^{+}$ is the correct output – a sentence different from $in$ but with the same chunk pattern. $out^{-}$ are $N_{negs}$ incorrect outputs, randomly chosen from the sentences that have a chunk pattern different from $in$. For each language, from the generated instances, we sample uniformly, based on the pattern of the input sentence, approximately 4000 instances, randomly split 80:20 into train:test. The train part is further split 80:20 into train:dev, resulting in a 2576:630:798 split for train:dev:test.
@@ 3 Probing the encoding of syntax
If the grammatical information in the sentences in our dataset, the sequences of chunks with specific properties relevant to the subject-verb agreement task (Figure 1 is an abstract form of knowledge within the pretrained model, it will be shared across languages), we would then see a high level of performance for a model trained on one of these languages, and tested on any of the other.
If however the grammatical information is a reflection of shallow language indicators, we would expect to see higher performance on languages that have overt grammatical number and chunk indicators, such as French and Italian, and a low rate of transfer of the information across languages.
@@
We use a two-level system illustrated in Figure 2, which separates detecting subject-verb agreement into two steps: (i) compress sentence embeddings into a representation that captures the sentence chunk structure and the relevant chunk properties (on the sentence level) (ii) use the compressed sentence representations to solve the BLM agreement problems, by detecting the pattern across the sequence of structures (on the task level). This architecture will allow us to test whether sentence structure – in terms of chunks – is shared across languages in a pretrained multilingual model.
Following previous explorations of sentence embedding properties [18, 11], all experiments use Electra [19] [^2] Electra pretrained model: google/electra-base-discriminator, and we take as sentence representations the embedding of the [CLS] token, reshaped as a two dimensional array with shape 32x24.
Data flow
An instance for a BLM problem consists of an ordered context sequence $S$ of sentences, $S=\{s_{i}|i=1,7\}$ as input, and an answer set $A$ with one correct answer $a_{c}$, and several incorrect answers $a_{err}$. Every sentence is embedded using the pretrained model. [^3] In the discussion that follows, when we say sentence we actually mean its embedding.
The two-level VAE system takes a BLM instance as input, decomposes its context sequence S into sentences and passes them individually as input to the sentence-level VAE. For each sentence $s_{i}\in S$, the system builds on-the-fly the candidate answers for the sentence level: the same sentence $s_{i}$ from input is used as the correct output, and a random selection of sentences from $S$ are the negative answers. After an instance is processed by the sentence level, for each sentence $s_{i}\in S$, we obtain its representation from the latent layer $l_{s_{i}}$, and reassemble the input sequence as $S_{l}=stack[l_{s_{i}}]$, and pass it as input to the task-level VAE. The loss function combines the losses on the two levels – a max-margin loss on the sentence level that contrasts the sentence reconstructed on the sentence level with the correct answer and the erroneous ones, and a max-margin loss on the task level that contrasts the answer constructed by the decoder with the answer set of the BLM instance.
Experiments
@@
Cross-lingual experiments, in which we train on data from one language, and then we test on all the others will provide information whether the patterns detected in sentence embeddings that encode chunk structure are transferable across languages or not.
The multilingual experiments, where we learn a model from data in all the languages, will provide additional clues – if the performance on testing on individual languages is higher than when training on that language alone, it means some information is shared across languages and can be beneficial.
@@ 4 Results
Figure 3 shows the results for the experiments on detecting chunk structure in sentence embeddings, in cross-lingual and multilingual training set-ups, for comparison (detailed results in Table 3).
@@
Two observations are relevant to our investigation: (i) while training and testing on the same language leads to good performance – indicating that the system does detect the chunk pattern in the input sentence embeddings – there is very little transfer effect. A slight effect is for the model learned on Italian, which has an improved performance, compared to the others, when applied to the French test data; (ii) learning using a multilingual training data leads to a deterioration of the performance, compared to learning in a monolingual setting. This again indicates that the information that is being learned, the chunk structure, is not shared across languages.
@@
An additional interesting insight comes from the analysis of the latent layer representations. Figure 4 shows the tSNE projection of the latent representations of the sentences in the training data.
Different colours show different chunk patterns, and different markers show different languages. Had the syntactic structure been shared, we had expected to see clusters for different patterns that cover all languages. Instead we note that each language seems to have its own pattern clusters.
@@
We note a similar phenomenon when the sentence structure detection is embedded in the system for solving the BLM agreement task, where an additional supervision signals comes from the task. Figure 5 shows the results for the multi-lingual and monolingual training set-ups for the type I data. Complete results are in Tables 4-5.
@@
Pretrained language models seem to acquire language understanding abilities from shallow cooccurrences through a lexical prediction task. The input information is transformed through several layers, various parts boosting each other through self-attention. Analysis of the architecture of transformer models, like BERT [4], have been able to localise and follow the flow of specific types of linguistic information through the system’s various layers [20, 3], to the degree that it seems as if the classical NLP pipeline is reflected in the succession of the model’s layers. Analysis of contextualized token embeddings shows that they can encode specific linguistic information, such as sentence structure [21], predicate argument structure [22], subjecthood and objecthood [23], among others. Sentence embeddings have also been probed using classifiers, and determined to encode specific types of linguistic information, such as subject-verb agreement [9], word order, tree depth, constituent information [24], auxiliaries[25] and argument structure [26]. It is not clear however how this information is encoded and accessed: does the pretrained model find deeper regularities and encodes abstract structures, or the predictions for all these tasks are based on shallow features that the probe used manages to assemble for the specific test it is used for? [27, 6]
We aimed to answer this question by using a multilingual set-up, and a simple syntactic structure detection task. The datasets we used – in 4 languages: English, French, Italian and Romanian –are approximately parallel from a lexical point of view, and parallel from a syntactic structure point of view. The property of interest is grammatical number, and the task is subject-verb agreement.
@@
The languages chosen share commonalities – French, Italian and Romanian are all Romance languages, English and French share much lexical material – but there are also differences: French and Italian use a similar manner to encode grammatical number. English has a very limited form of nominal plural morphology. In Romanian, number is expressed trough inflection, suffixation and case. These commonalities and differences help us interpret the results, and provide clues on how the targeted syntactic information is encoded.
@@
Previous experiments have shown that syntactic information – chunk sequences and their properties – can be accessed in transformer-based pretrained sentence embeddings [11]. In this multilingual set-up, we test whether this information has been identified based on language-specific shallow features, or whether the system has uncovered and encoded more abstract structures.
@@
The low rate of transfer for monolingual models and the decreased performance for a multi-lingually trained model for both our experimental set-ups indicate that the chunk sequence information is language specific and is assembled by the system based on shallow features. Further clues come from the fact that the only transfer happens between French and Italian, which encode phrases and grammatical number in a very similar manner. Embedding the sentence structure detection into a larger system, where it receives an additional learning signal (shared across languages) does not help to push towards finding a shared sentence representation space that encodes in a uniform manner the sentence structure shared across languages.

### v2

@@ Abstract
Questo lavoro chiede se i modelli linguistici multilingue preaddestrati catturino rappresentazioni linguistiche astratte valide attraverso svariate lingue. Il nostro approccio sviluppa dati sintetici curati su larga scala, con proprietà specifiche, e li utilizza per studiare le rappresentazioni di frasi costruite con modelli linguistici preaddestrati. Utilizziamo un nuovo task a scelta multipla e i dati afferenti, le Blackbird Language Matrices (BLM), per concentrarci su uno specifico fenomeno strutturale grammaticale - l’accordo tra il soggetto e il verbo - in diverse lingue. Per trovare la soluzione corretta a questo task è necessario un sistema che rilevi modelli e paradigmi linguistici complessi nelle rappresentazioni testuali. Utilizzando un’architettura a due livelli che risolve il problema in due fasi - prima impara gli oggetti sintattici e le loro proprietà nelle singole frasi e poi ne ricava gli elementi comuni - dimostriamo che, nonostante siano stati addestrati su testi multilingue in modo coerente, i modelli linguistici multilingue preaddestrati presentano differenze specifiche per ogni lingua e inoltre la struttura sintattica non è condivisa, nemmeno tra lingue tipologicamente molto vicine.
@@ 1 Introduction
To investigate this question, we use a seemingly simple task – subject-verb agreement. Subject-verb agreement is often used to test the syntactic abilities of deep neural networks [7, 8, 9, 10], because, while apparently simple and linear, it is in fact structurally, and theoretically, complex, and requires connecting the subject and the verb across arbitrarily long or complex structural distance. It has an added useful dimension – it relies on syntactic structure and grammatical number information that many languages share.
In previous work we have shown that simple structural information – the chunk structure of a sentence – which can be leveraged to determine subject-verb agreement, or to contribute towards more semantic tasks, can be detected in the sentence embeddings obtained from a pre-trained model [11]. This result, though, does not cast light on whether the discovered structure is deeper and more abstract, or it is rather just a reflection of surface indicators, such as function words or morphological markers.
To tease apart these two options, we set up an experiment covering four languages: English, French, Italian and Romanian. These languages, while different, have shared properties that make sharing of syntactic structure a reasonable expectation, if the pretrained multilingual model does indeed discover and encode syntactic structure. We use parallel datasets in the four languages, built by (approximately) translating the BLM-AgrF dataset [12], a multiple-choice linguistic test inspired from the Raven Progressive Matrices visual intelligence test, previously used to explore subject-verb agreement in French.
Our work offers two contributions: (i) four parallel datasets – on English, French, Italian and Romanian, focused on subject-verb agreement; (ii) cross-lingual and multilingual testing of a multilingual pretrained model, to explore the degree to which syntactic structure information is shared across different languages. Our cross-lingual and multilingual experiments show poor transfer across languages, even those most related, like Italian and French. This result indicates that pretrained models encode syntactic information based on shallow and language-specific clues, from which they are not yet able to take the step towards abstracting grammatical structure. The datasets are available at https://www.idiap.ch/dataset/(blm-agre|blm-agrf|blm-agri|blm_agrr) and the code at https://github.com/CLCL-Geneva/BLM-SNFDisentangling.
@@ 2 BLM task and BLM-Agr datasets
Inspired by existing IQ tests —Raven’s progressive matrices (RPMs)— we have developed a framework, called Blackbird Language Matrices (BLMs) [13] and several datasets [12, 14]. RPMs consist of a sequence of images, called the context, connected in a logical sequence by underlying generative rules [15]. The task is to determine the missing element in this visual sequence, the answer. The candidate answers are constructed to be similar enough that the solution can be found only if the rules are identified correctly.
@@
1 NP-sg PP1-sg VP-sg
2 NP-pl PP1-sg VP-pl
3 NP-sg PP1-pl VP-sg
4 NP-pl PP1-pl VP-pl
5 NP-sg PP1-sg PP2-sg VP-sg
6 NP-pl PP1-sg PP2-sg VP-pl
7 NP-sg PP1-pl PP2-sg VP-sg
@@
Answers
1 NP-pl PP1-pl PP2-sg VP-pl Correct
2 NP-pl PP1-pl et PP2-sg VP-pl Coord
3 NP-pl PP1-pl VP-pl WNA
4 NP-pl PP1-sg PP1-sg VP-pl WN1
5 NP-pl PP1-pl PP2-pl VP-pl WN2
6 NP-pl PP1-pl PP2-pl VP-sg AEV
7 NP-pl PP1-sg PP2-pl VP-sg AEN1
8 NP-pl PP1-pl PP2-sg VP-sg AEN2
Figure 1: BLM instances for verb-subject agreement, with two attractors. The errors can be grouped in two types: (i) sequence errors: WNA= wrong nr. of attractors; WN1= wrong gram. nr. for 1st attractor noun (N1); WN2= wrong gram. nr. for 2nd attractor noun (N2); (ii) grammatical errors: AEV=agreement error on the verb; AEN1=agreement error on N1; AEN2=agreement error on N2.
@@ 8 NP-pl PP1-pl PP2-sg VP-sg AEN2
Such an approach can be very useful for probing language models, as it allows to test whether they indeed detect the relevant linguistic objects and their properties, and whether (or to what degree) they use this information to find larger patterns. We have developed BLMs as a linguistic test. Figure 1 illustrates the template of a BLM subject-verb agreement matrix, with the different linguistic objects – chunks/phrases – and their relevant properties, in this case grammatical number. Examples in all languages under investigation are provided in Appendix B.
@@
A BLM problem for subject-verb agreement consists of a context set of seven sentences that share the subject-verb agreement phenomenon, but differ in other aspects – e.g. number of linearly intervening noun phrases between the subject and the verb (called attractors because they can interfere with the agreement), different grammatical numbers for these attractors, and different clause structures. The sequence is generated by a rule of progression of number of attractors, and alternation in the grammatical number of the different phrases. Each context is paired with a set of candidate answers generated from the correct answer by altering it to produce minimally contrastive error types. We have two types of errors (see Figure 1: (i) sequence errors – these candidate answers are grammatically correct, but they are not the correct continuation of the sequence; (ii) agreement errors – these candidate answers are grammatically erroneous, because the verb is in agreement with one of the intervening attractors. By constructing candidate answers with such specific error types, we can investigate the kind of information and structure learned.
The seed data for French was created by manually completing data previously published data [17]. From this initial data, we generated a dataset that comprises three subsets of increasing lexical complexity (details in [12]): Types I, II, III, corresponding to different amounts of lexical variation within a problem instance. Each subset contains three clause structures uniformly distributed within the data. The dataset used here is a variation of the BLM-AgrF [12] that separates sequence-based from other types of errors, to be able to perform deeper analyses into the behaviour of pretrained language models.
The datasets in English, Italian and Romanian were created by manually translating the seed French sentences into the other languages by native (Italian and Romanian) and near-native (English) speakers. The internal structure in these languages is very similar, so translations are approximately parallel. The differences lie in the treatment of preposition and determiner sequences that must be conflated into one word in some cases in Italian and French, but not in English. French and Italian use number-specific determiners and inflections, while Romanian and English encode grammatical number exclusively through inflections. In English most plural forms are marked by a suffix. Romanian has more variation, and noun inflections also encode case. Determiners are separate tokens, which are overt indicators of grammatical number and of phrase boundaries, whereas inflections may or may not be tokenized separately.
@@
The dataset consists of triples $(in,out^{+},Out^{-})$, where $in$ is an input sentence, $out^{+}$ is the correct output – a sentence different from $in$ but with the same chunk pattern. $Out^{-}$ are $N_{negs}=7$ incorrect outputs, randomly chosen from the sentences that have a chunk pattern different from $in$. For each language, we sample uniformly approx. 4000 instances from the generated data based on the pattern of the input sentence, randomly split 80:20 into train:test. The train part is split 80:20 into train:dev, resulting in a 2576:630:798 split for train:dev:test.
@@ 3 Probing the encoding of syntax
If the grammatical information in the sentences in our dataset – i.e. the sequences of chunks with specific properties relevant to the subject-verb agreement task (Figure 1) – is an abstract form of knowledge within the pretrained model, it will be shared across languages. We would then see a high level of performance for a model trained on one of these languages, and tested on any of the other. Additionally, when training on a dataset consisting of data in the four languages, the model should detect a shared parameter space that would lead to high results when testing on data for each language.
If however the grammatical information is a reflection of shallow language indicators, we expect to see higher performance on languages that have overt grammatical number and chunk indicators, such as French and Italian, and a low rate of cross-language transfer.
3.1 System architectures
A sentence-level VAE
To test whether chunk structure can be detected in sentence embeddings we use a VAE-like system, which encodes a sentence, and decodes a different sentence with the same chunk structure, using a set of contrastive negative examples – sentences that have different chunk structures from the input – to encourage the latent to encode the chunk structure.
The architecture of the sentence-level VAE is similar to a previously proposed system [18]: the encoder consists of a CNN layer with a 15x15 kernel, which is applied to a 32x24-shaped sentence embedding, followed by a linear layer that compresses the output of the CNN into a latent layer of size 5. The decoder mirrors the encoder.
An instance consists of a triple $(in,out^{+},Out^{-})$, where $in$ is an input sentence with embedding $e_{in}$ and chunk structure $p$, $out^{+}$ is a sentence with embedding $e_{out^{+}}$ with same chunk structure $p$, and $Out^{-}=\{s_{k}|k=1,N_{negs}\}$ is a set of $N_{negs}=7$ sentences with embeddings $e_{s_{k}}$, each with chunk pattern different from $p$ (and different from each other). The input $e_{in}$ is encoded into latent representation $z_{i}$, from which we sample a vector $\tilde{z}_{i}$, which is decoded into the output $\hat{e}_{in}$. To encourage the latent to encode the structure of the input sentence we use a max-margin loss function, to push for a higher similarity score for $\hat{e}_{in}$ with the sentence that has the same chunk pattern as the input ($e_{out^{+}}$) than the ones that do not. At prediction time, the sentence from the $\{out^{+}\}\cup Out^{-}$ options that has the highest score relative to the decoded answer is taken as correct.
@@ 3.1 System architectures
We use a two-level system illustrated in Figure 2, which separates the solving of the BLM task on subject-verb agreement into two steps: (i) compress sentence embeddings into a representation that captures the sentence chunk structure and the relevant chunk properties (on the sentence level) (ii) use the compressed sentence representations to solve the BLM agreement problems, by detecting the pattern across the sequence of structures (on the task level). This architecture will allow us to test whether sentence structure – in terms of chunks – is shared across languages in a pretrained multilingual model.
@@
All reported experiments use Electra [19] [^2] Electra pretrained model: google/electra-base-discriminator, with the sentence representations the embedding of the [CLS] token (details in [11]).
An instance for a BLM problem consists of an ordered context sequence $S$ of sentences, $S=\{s_{i}|i=1,7\}$ as input, and an answer set $A$ with one correct answer $a_{c}$, and several incorrect answers $a_{err}$. Every sentence is embedded using the pretrained model. To simplify the discussion, in the sections that follows, when we say sentence we actually mean its embedding.
The two-level VAE system takes a BLM instance as input, decomposes its context sequence $S$ into sentences and passes them individually as input to the sentence-level VAE. For each sentence $s_{i}\in S$, the system builds on-the-fly the candidate answers for the sentence level: the same sentence $s_{i}$ from input is used as the correct output, and a random selection of sentences from $S$ are the negative answers. After an instance is processed by the sentence level, for each sentence $s_{i}\in S$, we obtain its representation from the latent layer $l_{s_{i}}$, and reassemble the input sequence as $S_{l}=stack[l_{s_{i}}]$, and pass it as input to the task-level VAE. The loss function combines the losses on the two levels – a max-margin loss on the sentence level that contrasts the sentence reconstructed on the sentence level with the correct answer and the erroneous ones, and a max-margin loss on the task level that contrasts the answer constructed by the decoder with the answer set of the BLM instance (details in [11]).
3.2 Experiments
@@
Cross-lingual experiments – train on data from one language, test on all the others – show whether patterns detected in sentence embeddings that encode chunk structure are transferable across languages. The results on testing on the same language as the training provide support for the experimental set-up – the high results show that the pretrained language model used does encode the necessary information, and the system architecture is adequate to distill it.
The multilingual experiments, where we learn a model from data in all the languages, will provide additional clues – if the performance on testing on individual languages is comparable to when training on each language alone, it means some information is shared across languages and can be beneficial.
3.2.1 Syntactic structure in sentences
We use only the sentence level of the system illustrated in Figure 2 to explore chunk structure in sentences, using the data described in Section 2. For the cross-lingual experiments, the training dataset for each language is used to train a model that is then tested on each test set. For the multilingual setup, we assemble a common training data from the training data for all languages.
3.2.2 Solving the BLM agreement task
We solve the BLM agreement task using the two-level system, where a compacted sentence representation learned on the sentence level should help detect patterns in the input sequence of a BLM instance. Because the datasets are parallel, with shared sentence and sequence patterns, we test whether the added learning signal from the task level can help push the system to learn to map an input sentence into a representation that captures structure shared across languages. We perform cross-lingual experiments, where a model is trained on data from one language, and tested on all the test sets, and a multilingual experiment, where for each type I/II/III data, we assemble a training dataset from the training sets of the same type from the other languages. The model is then tested on the separate test sets.
3.3 Evaluation
For each training set we build three models, and plot the average F1 score. The standard deviation is very small, so we do not include it in the plot, but it is reported in the results Tables in Appendix C.
@@ 4 Results
Figure 3 shows the results for the experiments on detecting chunk structure in sentence embeddings, in cross-lingual and multilingual training setups, for comparison (detailed results in Table 3).
@@
Two observations are relevant to our investigation: (i) while training and testing on the same language leads to good performance – indicating that Electra sentence embeddings do contain relevant information about chunks, and that the system does detect the chunk pattern in these representations – there is very little transfer effect. A slight effect is detected for the model learned on Italian and tested on French; (ii) learning using multilingual training data leads to a deterioration of the performance, compared to learning in a monolingual setting. This again indicates that the system could not detect a shared parameter space for the information that is being learned, the chunk structure, and thus this information is encoded differently in the languages under study.
@@
An additional interesting insight comes from the analysis of the latent layer representations. Figure 4 shows the tSNE projection of the latent representations of the sentences in the training data after multilingual training. Different colours show different chunk patterns, and different markers show different languages. Had the information encoding syntactic structure been shared, the clusters for the same pattern in the different languages would overlap. Instead, we note that each language seems to have its own quite separate pattern clusters.
@@
When the sentence structure detection is embedded in the system for solving the BLM agreement task, where an additional supervision signals comes from the task, we note a similar result as when processing the sentences individually. Figure 5 shows the results for the multi-lingual and monolingual training setups for the type I data. Complete results are in Tables 4-5 in the appendix.
@@
Pretrained language models are learned from shallow cooccurrences through a lexical prediction task. The input information is transformed through several transformer layers, various parts boosting each other through self-attention. Analysis of the architecture of transformer models, like BERT [4], have localised and followed the flow of specific types of linguistic information through the system [20, 3], to the degree that the classical NLP pipeline seems to be reflected in the succession of the model’s layers. Analysis of contextualized token embeddings shows that they can encode specific linguistic information, such as sentence structure [21] (including in a multilingual set-up [22]), predicate argument structure [23], subjecthood and objecthood [24], among others. Sentence embeddings have also been probed using classifiers, and determined to encode specific types of linguistic information, such as subject-verb agreement [9], word order, tree depth, constituent information [25], auxiliaries[26] and argument structure [27].
Generative models like LLAMA seem to use English as the latent language in the middle layers [28], while other analyses of internal model parameters has lead to uncovering language agnostic and language specific networks of parameters [29], or neurons encoding cross-language number agreement information across several internal layers [30]. It has also been shown that subject-verb agreement information is not shared by BiLSTM models [31] or multilingual BERT [32]. Testing the degree to which word/sentence embeddings are multilingual has usually been done using a classification probe, for tasks like NER, POS tagging [33], language identification [34], or more complex tasks like question answering and sentence retrieval [35]. There are contradictory results on various cross-lingual model transfers, some of which can be explained by factors such as domain and size of training data, typological closeness of languages [36], or by the power of the classification probes. Generative or classification probes do not provide insights into whether the pretrained model finds deeper regularities and encodes abstract structures, or the predictions are based on shallower features that the probe used assembles for the specific test it is used for [37, 6].
@@
We aimed to answer this question by using a multilingual setup, and a simple syntactic structure detection task in an indirectly supervised setting. The datasets used – in English, French, Italian and Romanian – are (approximately) lexically parallel, and are parallel in syntactic structure. The property of interest is grammatical number, and the task is subject-verb agreement. The languages chosen share commonalities – French, Italian and Romanian are all Romance languages, English and French share much lexical material – but there are also differences: French and Italian use a similar manner to encode grammatical number, mainly through articles that can also signal phrase boundaries. English has a very limited form of nominal plural morphology, but determiners are useful for signaling phrase boundaries. In Romanian, number is expressed through inflection, suffixation and case, and articles are also often expressed through specific suffixes, thus overt phrase boundaries are less common than in French, Italian and English. These commonalities and differences help us interpret the results, and provide clues on how the targeted syntactic information is encoded.
@@
Previous experiments have shown that syntactic information – chunk sequences and their properties – can be accessed in transformer-based pretrained sentence embeddings [11]. In this multilingual setup, we test whether this information has been identified based on language-specific shallow features, or whether the system has uncovered and encoded more abstract structures.
@@
The low rate of transfer for the monolingual training setup and the decreased performance for the multilingual training setup for both our experimental configurations indicate that the chunk sequence information is language specific and is assembled by the system based on shallow features. Further clues come from the fact that the only transfer happens between French and Italian, which encode phrases and grammatical number in a very similar manner. Embedding the sentence structure detection into a larger system, where it receives an additional learning signal (shared across languages) does not help to push towards finding a shared sentence representation space that encodes in a uniform manner the sentence structure shared across languages.